OpenStudy (anonymous):

Find the quotient. 2x-3 7 ----- / ----- x x^2

OpenStudy (unklerhaukus):

is that \[\large \frac{2x-3}{x}\div\frac7{x^2}\]?

OpenStudy (anonymous):

yes @UnkleRhaukus

OpenStudy (anonymous):

@UnkleRhaukus hey can you reply to my message please

OpenStudy (unklerhaukus):

start by flipping the second fraction and changing the \(\div\) to \(\times\) \[=\large \frac{2x-3}{x}\times\frac{x^2}7\]

OpenStudy (unklerhaukus):

now you can cancel a factor of \(x\)

OpenStudy (unklerhaukus):

\[\dots\]

OpenStudy (anonymous):

is this the question you are working with?

OpenStudy (unklerhaukus):

@Claytonwj4